OpenStudy (anonymous):

what is the sum of an infinite geometric series with a first term of 6 and a common ratio of 1/2

OpenStudy (anonymous):

\[S_{\infty}=\frac{ a }{ 1-r }\]

OpenStudy (anonymous):

\[S _{\infty}=\frac{ 6 }{ 1-\frac{ 1 }{ 2} }\]

OpenStudy (anonymous):

ahh thank you!!